# Multimodal Fusion and Vision-Language Models: A Survey for Robot Vision

### v1

@@ Abstract
Robot vision has greatly benefited from advancements in multimodal fusion techniques and vision-language models (VLMs). We systematically review the applications of multimodal fusion in key robotic vision tasks, including semantic scene understanding, simultaneous localization and mapping (SLAM), 3D object detection, navigation and localization, and robot manipulation. We compare VLMs based on large language models (LLMs) with traditional multimodal fusion methods, analyzing their advantages, limitations, and synergies. Additionally, we conduct an in-depth analysis of commonly used datasets, evaluating their applicability and challenges in real-world robotic scenarios. Furthermore, we identify critical research challenges such as cross-modal alignment, efficient fusion strategies, real-time deployment, and domain adaptation, and propose future research directions, including self-supervised learning for robust multimodal representations, transformer-based fusion architectures, and scalable multimodal frameworks. Through a comprehensive review, comparative analysis, and forward-looking discussion, we provide a valuable reference for advancing multimodal perception and interaction in robotic vision. A comprehensive list of studies in this survey is available at https://github.com/Xiaofeng-Han-Res/MF-RV.
@@ 1 Introduction
Although several recent surveys have addressed multimodal fusion and VLMs, there remain significant limitations. Some surveys primarily focus on traditional perception tasks such as semantic segmentation [27] or 3D object detection [28], lacking a systematic review of deep learning-driven cross-modal modeling approaches. Most of them do not cover key topics such as SLAM, embodied navigation, multimodal 3D detection, or visual localization.On the other hand, recent surveys on VLMs tend to emphasize model architectures and natural language processing tasks, but fall short of offering a comprehensive discussion on their applications in robot perception and control. For example, Li et al. [29] presented an evolutionary overview of mainstream VLM architectures, but did not explore their use in tasks like navigation or robotic manipulation. The introductory survey by Bordes et al. [30] covers training paradigms and evaluation methods, but is mainly intended as a beginner-friendly guide and fails to address the core challenges of multimodal fusion in robotic vision. In addition, although Wu et al. [31] provide an in-depth analysis of multimodal techniques for embodied navigation, their work does not include emerging vision-language model architectures, nor does it place tasks such as SLAM and manipulation within a unified analytical framework.In contrast to the aforementioned works, this survey focuses on the systematic integration and comparison of multimodal fusion and vision-language models within the context of robot vision tasks, covering key applications ranging from semantic scene understanding to embodied navigation, SLAM, and robotic manipulation. We not only analyze and compare traditional multimodal fusion methods with emerging foundation models, but also evaluate their applicability to real-world robotic systems, computational cost, and potential for future scalability. In addition, this survey highlights often overlooked research directions such as cross-modal self-supervised learning and lightweight fusion architectures, with the aim of providing forward-looking insights for the development of more intelligent, generalizable, and deployable robot vision systems.
@@ 2.1 Semantic Scene Understanding
Figure 3: The above diagram intuitively illustrates the basic processes of three main strategies for multimodal fusion: early fusion, mid fusion, and late fusion. In the figure, each modality data (such as modality data1, modality data2, etc.) is independently processed by a feature extractor. Early fusion directly fuses data from different modalities before feature extraction; Mid term fusion combines modal features through specific mechanisms such as feature concatenation or weighting after extracting them; Late stage fusion is achieved by integrating the decision results of each modality after independent decision-making is completed. The diagram clearly reflects the key differences and roles of the three fusion methods in the multimodal processing flow.
@@ 2.1.1 Encoder-Decoder Framework
Early multimodal fusion methods primarily relied on the encoder-decoder framework to extract and combine information from different modalities. However, with the advancement of research, more recent approaches have integrated various attention mechanisms with the encoder-decoder framework to enhance the flexibility and accuracy of information fusion. These attention mechanisms effectively guide the model to focus on key features across modalities, thereby improving the performance of multimodal learning, especially in handling complex tasks.
@@ 2.1.2 Attention Mechanism
Several recent studies have exemplified the effectiveness of attention-based frameworks in multimodal scene understanding, particularly in 3D scene completion, segmentation, and object detection. For instance,MRFTrans [4], a transformer-based framework, fuses semantic, geometric, and depth information to address the challenge of monocular 3D scene completion, achieving state-of-the-art performance on SemanticKITTI and NYUv2 datasets by effectively modeling long-range dependencies and refining depth estimates. DefFusion [14] introduces a deformable transformer mechanism that dynamically adjusts to multimodal data, mitigating over-attention and noise, and significantly improving 3D semantic segmentation, particularly on large-scale benchmarks like NuScenes.On the other hand ,DMA [59] focuses on open-vocabulary 3D scene understanding, aligning dense correspondences between 3D points, 2D images, and text, utilizing vision-language models to generalize across diverse indoor and outdoor scenarios. Lastly, Multi-Sem Fusion [60] integrates LiDAR and camera data for 3D object detection, employing adaptive attention and deep feature fusion to enhance detection accuracy, although its reliance on offline processing limits real-time applicability. The CLFT [61] model achieves the fusion of camera and LiDAR data by projecting LiDAR point clouds onto the camera plane, upsampling them into 2D feature maps, and then inputting them together with camera images into a Transformer-based network. The model utilizes a bidirectional network and a cross-fusion strategy to integrate data from both sensors at the Transformer decoder layer, enhancing the accuracy and robustness of semantic segmentation in autonomous driving.
@@ 2.1.3 Graph Neural Network
Figure 4: The above diagram intuitively illustrates the workflow of multimodal fusion using Graph Neural Networks (GNN). In the figure, each modality data (such as modality data1, modality data2, etc.) is first transformed into a graph structure through a graph construction process. These graphs are then processed by separate GNNs to extract high-level representations. The final decision stage integrates the outputs from different GNNs, enabling effective multimodal reasoning. This diagram clearly demonstrates the role of GNNs in learning structured relationships within each modality and highlights the fusion process in multimodal understanding.
@@
Table 1: Representative multimodal fusion methods categorized by fusion strategy.
Multimodal fusion Type Representative Methods
Encoder-Decoder Framework
DeepLabv3+ [54],
PointRend [55],
HRNet [56],
DDRNet [57],
MAMT [58],
PIDNet [12],
Attention Mechanism
MRFTrans [4],
DefFusion [14],
DMA [59],
Multi-Sem Fusion [60],
CLFT [61],
OccFusion [64],
SENET [63],
CONet [62],
HyDRa [16],
Graph Neural Network
MISSIONGNN [65],
VQA-GNN [66],
MuSe-GNN [67],
CCA-GNN [68],
MGA [69],
MGNNS [70],
@@ 2.2.1 Lidar+Camera
3D object detection is one of the crucial perception tasks in autonomous driving systems, whose primary goal is to obtain three-dimensional information about the environment through various sensors, identify and locate targets such as vehicles, pedestrians, and obstacles, and then provide support for path planning and decision-making. In autonomous driving, common sensors include Cameras, LiDARs, Radars, and ultrasonic sensors, etc. A single sensor often cannot meet the requirements for safe and reliable perception. Multi-modal fusion combines the advantages of different types of sensors to make up for their respective shortcomings, thus providing more comprehensive and accurate environmental perception capabilities.
Camera-based single-modal 3D object detection [1, 72, 2, 3] infers 3D information from 2D images. While RGB images provide rich visual data, the lack of direct depth perception requires complex estimation techniques, making detection vulnerable to occlusions, poor angles, and adverse weather conditions. These factors degrade accuracy and robustness. In contrast, LiDAR-based detection [73, 74, 75, 76, 77] processes 3D point cloud data, offering precise depth estimation and spatial structure representation. However, LiDAR data is often sparse, especially at long distances or on low-reflectivity surfaces, making it challenging to detect small or distant objects. Additionally, the absence of color and texture limits its ability to distinguish visually similar but categorically different objects. Therefore, although single-modal 3D object detection methods each have their own advantages, they all have certain limitations. To solve this problem, more and more multi-modal fusion strategies are adopted in autonomous driving systems, among which the fusion of Cameras and LiDAR point clouds is a main research direction.
@@
Attention-based methods enhance object detection by dynamically focusing on key multimodal information. 3D-CVF [81] employs adaptive gating fusion with attention maps for feature weighting. FusionPainting [83] integrates 2D and 3D segmentation to address boundary blurring. DeepFusion [85] ensures precise alignment of LiDAR and image features using InverseAug and cross-attention mechanisms. Unlike hard association methods [5, 71, 81], TransFusion [13] introduces a soft association fusion framework using a Transformer decoder for high-precision 3D detection. The first layer generates initial 3D boxes with object queries, while the second layer refines them using cross-attention. Additionally, an image-guided query strategy improves detection in sparse point clouds. UVTR [84] converts images into voxel representations, enabling cross-modal learning. DeepInteraction [86] addresses the limitation of unilateral fusion by introducing a multimodal interaction encoder and decoder. CMT [87] enhances object localization by encoding 3D coordinates and using a Transformer decoder for multimodal interaction. BEVFusion [88] aligns image and point cloud features in BEV space using a convolutional encoder. However, its camera features lack spatial awareness, which GAFusion [89] improves using Sparse Depth Guidance and Laser Occupancy Guidance. Additionally, it employs a Laser-Guided Adaptive Fusion Transformer and a temporal fusion module for multi-frame aggregation. IS-Fusion [90] captures both instance-level and scene-level context using Hierarchical Scene Fusion and Instance-Guided Fusion, effectively integrating multimodal information at different granularities.
@@
As shown in Figure 6, in the field of LiDAR-Camera fusion, we have witnessed the evolution from early simple feature level fusion to today’s complex attention mechanisms and Transformer architectures. These advanced fusion strategies have not only improved the accuracy of 3D object detection but also enhanced the model’s adaptability and robustness to complex environments. With the advancement of deep learning and sensor technology, future fusion methods are expected to achieve higher precision and lower computational costs, while also better handling challenges in dynamic scenes, such as occlusions and rapidly moving objects.
@@ 2.3.1 Embodied Navigation
The construction of topological maps plays a crucial role in enhancing robotic navigation capabilities. As shown in Figure 8, navigation agents can dynamically update environmental models, integrating current observations with existing topological information to progressively expand navigable regions. Compared to relying solely on local decision-making, incorporating topological maps allows robots to optimize navigation paths based on global information, effectively mitigating random exploration-induced disorientation. Furthermore, when navigation errors occur, robots can trace back historical trajectories and dynamically adjust paths. For instance, DUET [115] employs a dual-scale graph transformer, integrating local visual features with global spatial information during topological path construction to achieve cross-modal alignment and optimize navigation strategies. BEVBert [116] further enhances pre-training on hybrid topo-metric maps, enabling robots to better comprehend spatial structures and improve cross-modal reasoning capabilities in complex environments. Chasing Ghosts [117] introduces Bayesian state tracking, integrating semantic spatial maps to enhance trajectory inference and goal localization, thereby improving navigation robustness in uncertain environments.Additionally, long-horizon planning is essential for embodied navigation. PREVALENT [118] and PRET [119] leverage Vision-Language Navigation (VLN) pre-training and directed fidelity trajectories to enhance alignment between visual and linguistic information, thereby optimizing path search efficiency. HAMT [120] introduces a History-Aware Multimodal Transformer (HAMT), explicitly encoding comprehensive navigation history during cross-modal decision-making, reinforcing navigation capabilities in long-horizon tasks. Meanwhile, TD-STP [121] employs a Target-Driven Structured Transformer Planner (TD-STP), integrating topological maps to predict long-term navigation goals and modeling environmental topologies, ensuring efficient adaptation and optimized action planning for robotic navigation. NavGPT-2 [10] combines large-scale Vision-Language Models (LLMs), utilizing frozen LLMs aligned with visual content to enhance navigation reasoning, while maintaining explainability and human-interaction capabilities.
@@
These methods, such as InstructNav [9], NaVid [114], BEVBert [116], and CVDN [124], demonstrate the transformative potential of multimodal, fusion in embodied navigation. By integrating visual input, language processing, and spatial reasoning, robots can now follow natural language instructions, navigate toward visual goals, or perform goal-directed navigation tasks with greater flexibility and adaptability [114, 126, 127, 128]. The fusion of different sensory modalities such as vision, language, and spatial cues enables robots to navigate more robustly in diverse and dynamic environments, pushing the boundaries of autonomous navigation systems.

### v3

@@ Abstract
Robot vision has greatly benefited from advancements in multimodal fusion techniques and vision-language models (VLMs). We adopt a task-oriented perspective to systematically review the applications and advancements of multimodal fusion methods and VLMs in the field of robot vision. For semantic scene understanding tasks, we categorize fusion approaches into encoder-decoder frameworks, attention-based architectures, and graph neural networks. Meanwhile, we also analyze the architectural characteristics and practical implementations of these fusion strategies in key tasks such as simultaneous localization and mapping (SLAM), 3D object detection, navigation, and manipulation. We compare the evolutionary paths and applicability of VLMs based on large language models (LLMs) with traditional multimodal fusion methods.Additionally, we conduct an in-depth analysis of commonly used datasets, evaluating their applicability and challenges in real-world robotic scenarios. Building on this analysis, we identify key challenges in current research, including cross-modal alignment, efficient fusion, real-time deployment, and domain adaptation. We propose future directions such as self-supervised learning for robust multimodal representations, structured spatial memory and environment modeling to enhance spatial intelligence, and the integration of adversarial robustness and human feedback mechanisms to enable ethically aligned system deployment. Through a comprehensive review, comparative analysis, and forward-looking discussion, we provide a valuable reference for advancing multimodal perception and interaction in robotic vision. A comprehensive list of studies in this survey is available at https://github.com/Xiaofeng-Han-Res/MF-RV.
@@ 1 Introduction
Table 1: Comparison of existing surveys and ours in terms of task scope and technical focus. Tasks: Number of covered tasks, including semantic scene understanding, SLAM, 3D object detection, navigation and localization, and robot manipulation. Arch.: Vision-Language Model architecture analysis; CM-SSL: Cross-modal self-supervised learning ;LightFusion: Lightweight fusion methods.
Survey #Tasks Arch. CM-SSL LightFusion
Zhang et al. [27] 2
Alaba et al. [28] 1
Li et al. [29] 1
Wu et al. [30] 1
Bordes et al. [31] 1
Lin et al. [32] 1
Chen et al. [33] 1
Abdulmaksoud et al. [34] 2
Ours 5
Although recent years have seen several surveys on multimodal fusion and vision-language models, significant limitations remain in terms of task coverage and modeling depth (Table 1). On one hand, some surveys, such as Zhang et al.[27] and Alaba et al.[28], focus on specific traditional tasks (e.g., semantic segmentation, 3D detection), lacking systematic coverage of critical robotic tasks such as SLAM, embodied navigation, multimodal 3D perception, and manipulation. On the other hand, studies like Li et al.[29] and Bordes et al.[31] explore the architectural evolution and evaluation paradigms of VLMs, but their analyses are primarily centered on language generation tasks, with insufficient discussion on the practical deployment of vision-language models in real robotic systems.Moreover, although Wu et al.[30] and Lin et al.[32] focus on the integration of perception and decision-making in embodied navigation, they do not incorporate tasks such as SLAM and manipulation into a unified framework. Meanwhile, Abdulmaksoud et al.[34] and Chen et al.[33] provide systematic reviews of sensor fusion methods and LiDAR-Vision SLAM, respectively, yet they do not explore the potential integration of VLMs with self-supervised alignment strategies. In contrast, this survey focuses on the systematic integration and comparison of multimodal fusion methods and vision-language models (VLMs) in the context of robotic vision tasks, covering key applications ranging from semantic scene understanding[35] to embodied navigation, SLAM, and robotic manipulation. Moreover, we not only systematically analyze and compare traditional multimodal fusion methods and VLMs, but also highlight often-overlooked research directions such as cross-modal self-supervised learning and lightweight fusion architectures. Our goal is to provide forward-looking insights for building more intelligent, generalizable, and deployable robotic vision systems.
@@ 2.1 Semantic Scene Understanding
Figure 3: The above diagram intuitively illustrates the basic processes of three main strategies for multimodal fusion: early fusion, mid fusion, and late fusion. In the figure, each modality data (such as images, audio, Lidar, text, etc.) is independently processed by a feature extractor. Early fusion directly fuses data from different modalities before feature extraction; Mid term fusion combines modal features through specific mechanisms such as feature concatenation or weighting after extracting them; Late stage fusion is achieved by integrating the decision results of each modality after independent decision-making is completed. The diagram clearly reflects the key differences and roles of the three fusion methods in the multimodal processing flow.
@@ 2.1.1 Encoder-Decoder Framework
Early multimodal fusion methods primarily relied on the encoder-decoder framework to extract and combine information from different modalities. Nevertheless, as research has progressed, contemporary methodologies have incorporated a range of attention mechanisms alongside the encoder-decoder framework, thereby augmenting the flexibility and precision of information fusion. These attention mechanisms effectively guide the model to focus on key features across modalities, thereby improving the performance of multimodal learning, especially in handling complex tasks.Recent innovations such as elastic deep multi-view autoencoders with diversity embedding [63] further enhance the diversity and robustness of multimodal representations, complementing attention-based fusion strategies.
@@ 2.1.2 Attention Mechanism
Figure 4: Illustration of the standard self-attention mechanism. The query vector is compared against all key vectors using a compatibility function $f(Q,K)$, typically dot product. The resulting scores $s_{1},s_{2},s_{3}$ are normalized via the Softmax function to obtain attention weights $a_{1},a_{2},a_{3}$, which are then used to compute a weighted sum over the value vectors $V_{1},V_{2},V_{3}$, producing the final attention output.
Several recent studies have exemplified the effectiveness of attention-based frameworks in multimodal scene understanding, particularly in 3D scene completion, segmentation, and object detection. For instance,MRFTrans [4], a transformer-based framework, fuses semantic, geometric, and depth information to address the challenge of monocular 3D scene completion, achieving state-of-the-art performance on SemanticKITTI and NYUv2 datasets by effectively modeling long-range dependencies and refining depth estimates. DefFusion [14] introduces a deformable transformer mechanism that dynamically adjusts to multimodal data, mitigating over-attention and noise, and significantly improving 3D semantic segmentation, particularly on large-scale benchmarks like NuScenes.On the other hand ,DMA [64] focuses on open-vocabulary 3D scene understanding, aligning dense correspondences between 3D points, 2D images, and text, utilizing vision-language models to generalize across diverse indoor and outdoor scenarios. Lastly, Multi-Sem Fusion [65] integrates LiDAR and camera data for 3D object detection, employing adaptive attention and deep feature fusion to enhance detection accuracy, although its reliance on offline processing limits real-time applicability. The CLFT [66] model achieves the fusion of camera and LiDAR data by projecting LiDAR point clouds onto the camera plane, upsampling them into 2D feature maps, and then inputting them together with camera images into a Transformer-based network. The model utilizes a bidirectional network and a cross-fusion strategy to integrate data from both sensors at the Transformer decoder layer, enhancing the accuracy and robustness of semantic segmentation in autonomous driving. However, it is important to note that although the attention mechanism performs well in multimodal alignment, the standard self-attention operation is computationally expensive, with its time and memory complexity growing quadratically with input length,as show in the standard self-attention operation is defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,$ (1)
as illustrated in Figure 4, where $Q,K,V\in\mathbb{R}^{L\times d_{k}}$ are the query, key, and value matrices, $L$ is the input sequence length, and $d_{k}$ is the feature dimension. This operation requires pairwise interactions between all tokens in the sequence. Its computational complexity is:
$\mathcal{O}(L^{2}\cdot d_{k}),$ (2)
which becomes a bottleneck for long sequences or high-resolution visual inputs in resource-constrained environments.This challenge is further intensified in multimodal settings where high-resolution images and long text sequences must be processed jointly. Although approaches such as sparse attention, linear attention, and low-rank approximations have been proposed to alleviate this burden, a trade-off still exists between efficiency and alignment accuracy, which must be carefully balanced in time-critical applications.
@@ 2.1.3 Graph Neural Network
Figure 5: The above diagram intuitively illustrates the workflow of multimodal fusion using Graph Neural Networks (GNN). In the figure, each modality data (such as images, audio, Lidar, text, etc.) is first transformed into a graph structure through a graph construction process. These graphs are then processed by separate GNNs to extract high-level representations. The final decision stage integrates the outputs from different GNNs, enabling effective multimodal reasoning. This diagram clearly demonstrates the role of GNNs in learning structured relationships within each modality and highlights the fusion process in multimodal understanding.
@@
Table 2: Representative multimodal fusion methods categorized by fusion strategy and publication year.
Multimodal fusion Type Representative Methods Year
Encoder-Decoder Framework PIDNet [12] 2023
MAMT [62] 2022
DDRNet [61] 2021
PointRend [59] 2020
HRNet [60] 2019
DeepLabv3+ [58] 2018
Attention Mechanism MRFTrans [4] 2024
DefFusion [14] 2024
DMA [64] 2024
Multi-Sem Fusion [65] 2024
CLFT [66] 2024
OccFusion [69] 2024
HyDRa [16] 2024
CONet [67] 2023
SENET [68] 2018
Graph Neural Network MISSIONGNN [70] 2024
VQA-GNN [71] 2023
MuSe-GNN [72] 2023
CCA-GNN [73] 2023
MGA [74] 2023
MGNNS [75] 2021
@@ 2.2.1 Lidar+Camera
3D object detection is a crucial perception task in autonomous driving systems. Its primary goal is to acquire three-dimensional environmental information through various sensors, identify and localize targets (e.g., vehicles, pedestrians, obstacles), and provide support for path planning and decision-making.In autonomous driving, common sensors include Cameras, LiDARs, Radars, and ultrasonic sensors, etc. A single sensor often cannot meet the requirements for safe and reliable perception. Multi-modal fusion combines the advantages of different types of sensors to make up for their respective shortcomings, thus providing more comprehensive and accurate environmental perception capabilities.
Camera-based single-modal 3D object detection [1, 77, 2, 3] infers 3D information from 2D images. While RGB images provide rich visual data, the lack of direct depth perception requires complex estimation techniques, making detection vulnerable to occlusions, poor angles, and adverse weather conditions. These factors degrade accuracy and robustness. In contrast, LiDAR-based detection [78, 79, 80, 81] processes 3D point cloud data, offering precise depth estimation and spatial structure representation. However, LiDAR data is often sparse, especially at long distances or on low-reflectivity surfaces, making it challenging to detect small or distant objects. Additionally, the absence of color and texture limits its ability to distinguish visually similar but categorically different objects. Therefore, although single-modal 3D object detection methods each have their own advantages, they all have certain limitations. To solve this problem, more and more multi-modal fusion strategies are adopted in autonomous driving systems, among which the fusion of Cameras and LiDAR point clouds is a main research direction.
@@
Attention-based methods enhance object detection by dynamically focusing on key multimodal information. 3D-CVF [85] employs adaptive gating fusion with attention maps for feature weighting. FusionPainting [87] integrates 2D and 3D segmentation to address boundary blurring. DeepFusion [89] ensures precise alignment of LiDAR and image features using InverseAug and cross-attention mechanisms. Unlike hard association methods [5, 76, 85], TransFusion [13] introduces a soft association fusion framework using a Transformer decoder for high-precision 3D detection. The first layer generates initial 3D boxes with object queries, while the second layer refines them using cross-attention. Additionally, an image-guided query strategy improves detection in sparse point clouds. UVTR [88] converts images into voxel representations, enabling cross-modal learning. DeepInteraction [90] addresses the limitation of unilateral fusion by introducing a multimodal interaction encoder and decoder. CMT [91] enhances object localization by encoding 3D coordinates and using a Transformer decoder for multimodal interaction. BEVFusion[92] aligns image and point cloud features in a unified BEV space, achieving high accuracy and efficiency in multimodal 3D perception tasks, particularly in large-scale scenarios such as autonomous driving. However, its image branch lacks spatial geometric modeling capabilities, making it prone to spatial reasoning errors in scenarios involving occlusions, viewpoint variations, or complex structures. In robotic vision tasks, although BEV representations provide a comprehensive global view, they are susceptible to degraded robustness in unstructured environments due to sparse point clouds, image blur, and depth estimation noise. Furthermore, the method relies on multi-modal projection and fusion operations that are computationally expensive, posing challenges for real-time deployment on resource-constrained platforms. The follow-up work GAFusion[93] addresses the spatial awareness limitations by introducing sparse depth guidance and laser occupancy maps, and incorporates a laser-guided adaptive fusion transformer along with a temporal fusion module to enhance inter-modal and inter-frame consistency. However, these enhancements further increase the model complexity and inference cost, limiting its practicality on embedded robotic systems, especially in scenarios requiring both high frame rates and low power consumption.IS-Fusion [94] employs instance-guided fusion and hierarchical scene modeling to jointly capture fine-grained semantics and global contextual information, demonstrating strong performance in object detection and 3D scene understanding. While the method excels in structured and object-dense autonomous driving environments, it faces limitations in robotic applications involving dynamic, cluttered, or partially observable scenes, where its reliance on high-quality instance features can degrade performance. For example, in conditions with occlusions, low lighting, or incomplete observations, instance masks may be difficult to obtain accurately, reducing the effectiveness of the fusion process. Moreover, the method’s multi-level fusion architecture introduces substantial computational overhead and lower inference efficiency, posing challenges for deployment in control-critical robotic systems. Additionally, as its experiments are primarily based on autonomous driving datasets, the method’s generalization to common indoor tasks such as navigation and manipulation in service robots remains to be further validated.
@@
As shown in Figure 7, in the field of LiDAR-Camera fusion, we have witnessed the evolution from early simple feature level fusion to today’s complex attention mechanisms and Transformer architectures. As shown in Table 3, Transformer architectures have gradually become mainstream, giving rise to a number of new models such as TransFusion, BEVFusion, and MV2DFusion. By leveraging multi-head attention mechanisms and modality-specific structures, these models enable more flexible and efficient feature fusion, enhancing the model’s ability to understand complex scenes and multimodal semantics.These advanced fusion strategies have not only improved the accuracy of 3D object detection but also enhanced the model’s adaptability and robustness to complex environments. With the advancement of deep learning and sensor technology, future fusion methods are expected to achieve higher precision and lower computational costs, while also better handling challenges in dynamic scenes, such as occlusions and rapidly moving objects.
@@ 2.3.1 Embodied Navigation
The construction of topological maps plays a crucial role in enhancing robotic navigation capabilities. As shown in Figure 9, navigation agents can dynamically update environmental models, integrating current observations with existing topological information to progressively expand navigable regions. Compared to relying solely on local decision-making, incorporating topological maps allows robots to optimize navigation paths based on global information, effectively mitigating random exploration-induced disorientation. Furthermore, when navigation errors occur, robots can trace back historical trajectories and dynamically adjust paths. For instance, DUET [119] employs a dual-scale graph transformer, integrating local visual features with global spatial information during topological path construction to achieve cross-modal alignment and optimize navigation strategies. BEVBert [120] further enhances pre-training on hybrid topo-metric maps, enabling robots to better comprehend spatial structures and improve cross-modal reasoning capabilities in complex environments. Chasing Ghosts [121] introduces Bayesian state tracking, integrating semantic spatial maps to enhance trajectory inference and goal localization, thereby improving navigation robustness in uncertain environments.Additionally, long-horizon planning is essential for embodied navigation. PREVALENT [122] and PRET [123] leverage Vision-Language Navigation (VLN) pre-training and directed fidelity trajectories to enhance alignment between visual and linguistic information, thereby optimizing path search efficiency. HAMT [124] introduces a History-Aware Multimodal Transformer (HAMT), explicitly encoding comprehensive navigation history during cross-modal decision-making, reinforcing navigation capabilities in long-horizon tasks. Meanwhile, TD-STP [125] employs a Target-Driven Structured Transformer Planner (TD-STP), integrating topological maps to predict long-term navigation goals and modeling environmental topologies, ensuring efficient adaptation and optimized action planning for robotic navigation. PRET [123] proposes a directionally faithful trajectory planning strategy that enhances planning efficiency and significantly reduces computational overhead by evaluating the semantic alignment between language instructions and candidate paths to select the most suitable goal node. It employs a direction-aware graph structure to avoid frequent updates to the entire map, making it particularly suitable for structured indoor environments. However, its deployment on real-world robotic platforms still faces multiple challenges. The dynamic nature of real environments, noisy sensory inputs, and reduced accuracy in trajectory-instruction alignment can all negatively affect navigation performance. Moreover, PRET requires continuous maintenance of global path information, imposing high demands on real-time perception synchronization and computational resources. NavGPT-2 [10], on the other hand, integrates large-scale vision-language models (e.g., InstructBLIP) with a topological path planning module, adopting a decoupled design that separates language reasoning from navigation policy learning. This allows for both semantic interpretability and efficient decision-making. By freezing the large model and only training a lightweight policy network, NavGPT-2 achieves strong performance in instruction comprehension, path visualization, and human-robot interaction, demonstrating good data efficiency and generalization capability. However, it still relies on the pretrained model’s visual-semantic alignment ability, which may degrade under real-world conditions such as low lighting, occlusion, or sensory noise. Furthermore, the model’s long reasoning pipeline and complex prompt design present efficiency bottlenecks when deployed on resource-constrained embedded platforms.
@@
These methods, such as InstructNav [9], NaVid [118], BEVBert [120], and CVDN [128], demonstrate the transformative potential of multimodal, fusion in embodied navigation. By integrating visual input, language processing, and spatial reasoning, robots can now follow natural language instructions, navigate toward visual goals, or perform goal-directed navigation tasks with greater flexibility and adaptability [118, 130, 131, 132, 133]. The fusion of different sensory modalities such as vision, language, and spatial cues enables robots to navigate more robustly in diverse and dynamic environments, pushing the boundaries of autonomous navigation systems.